OpenStudy (anonymous):

Write the given sentence as an equation. Tim's age in 7 years will be three times what it was 19 years ago.

OpenStudy (anonymous):

Make whatever you don't know a variable. You don't know tim's age 7 years ago and you don't know tim's age 19 years ago. Tim's age in 7 years = x Tim's age 19 years ago = y (Tim's age in 7 years) (will be) (three times) (what it was 19 years ago.) x = 3 * y so x=3y

OpenStudy (anonymous):

But the choises are 3(t+7)=t-9 and t+19=3(t-7) and t+7=3(t-19) and 3(t+19)=1-7